# Character Theory for Semilinear Representations

### v2

@@ 9. Semilinear Schur Index
For $W\in\operatorname{Irr}_{L}(H)$, set $m_{K}^{L}(W)\coloneqq m(V)$ for the unique corresponding $V\in\operatorname{Irr}_{L}^{\rtimes}(G)$.
@@
Let $W\in\operatorname{Irr}_{L}(H)$. Then $W$ is extendable to a semilinear representation of $G$ if and only if $\Gamma$ fixes $W$ and $m_{K}^{L}(W)=1$.
From the results of the previous section, the numbers $m_{K}^{L}(W)$ have the following properties. Recall that we write $L_{W}\coloneqq L^{\Gamma_{W}}$, and that $G_{W}$ is the preimage of $\Gamma_{W}$ in $G$. For any intermediate field $E$ with $K\leq E\leq L$, we write $\Gamma_{E}$ for the preimage of $\operatorname{Gal}(L/E)$ under the isomorphism $\Gamma\xrightarrow{\sim}\operatorname{Gal}(L/K)$.
@@
Suppose that $W\in\operatorname{Irr}_{L}(H)$, and set $\mathcal{O}(W)\coloneqq\bigoplus_{\gamma\in\Gamma/\Gamma_{W}}\gamma*W$. Then:
@@
    $\mathcal{O}(W)^{\oplus m_{K}^{L}(W)}=V|_{H}$ for a unique $V\in\operatorname{Irr}_{L}^{\rtimes}(G)$,
@@
    $m_{K}^{L}(W)$ is the unique $m\geq 1$ such that $\mathcal{O}(W)^{\oplus m}$ extends to an irreducible object of $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$,
@@
    For any semisimple $V\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$, the multiplicity of $W$ in $V|_{H}$ is a multiple of $m_{K}^{L}(W)$.
@@
    $W^{\oplus m_{L_{W}}^{L}(W)}=U|_{H}$ for a unique $U\in\operatorname{Irr}_{L}^{\rtimes}(G_{W})$,
@@
    $m_{L_{W}}^{L}(W)$ is the unique $m\geq 1$ such that $W^{\oplus m}$ extends to an irreducible object of $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G_{W})$,
@@
    For any semisimple $U\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G_{W})$, the multiplicity of $W$ in $U|_{H}$ is a multiple of $m_{L_{W}}^{L}(W)$.
@@
Points (1), (2), (4), (5) and (6) follow from Theorem 8.20 and Corollary 8.6. For point (3), if $V|_{H}=\mathcal{O}(W)^{\oplus m}$ for some $m\geq 1$, then $V$ is semisimple by Corollary 8.14, and the claim follows from (4). For point (7), we may take $S$ to be $G_{W}$ in Proposition 8.23, to obtain that $m_{L_{W}}^{L}(W)\mid m_{K}^{L}(W)$ and $m_{K}^{L}(W)\mid m_{L_{W}}^{L}(W)\cdot[\Gamma_{W}:(\Gamma_{W})_{W}]=m_{L_{W}}^{L}(W)$.
@@
with kernel $C_{n}=\langle x\rangle\subset C_{np}=\langle y\rangle$, where $y^{p}=x$. Then for any $1$-dimensional representation $W$ of $H$ valued in $K$, then similarly to above, $m_{K}^{L}(W)=1$ or $m_{K}^{L}(W)=p$, and $m_{K}^{L}(W)=1$ if and only if
$N_{K}^{L}(a)=\chi_{W}(x)$
for some $a\in L^{\times}$.
@@ 10. Recovery of the Classical Schur Index
Let $V\in\operatorname{Irr}_{L}^{\rtimes}(G)$ and $W\in\operatorname{Irr}_{L}(H)$ where $W$ is an irreducible submodule of $V_{L}$, or equivalently that $V$ is an irreducible submodule of $W|_{K[H]}$. Then:
@@
    for some integer $\overline{m}_{K}^{L}(W)\geq 1$.
  • In particular, $V_{L}$ and $W|_{K[H]}$ are semisimple, and $\overline{m}_{K}^{L}(W)\mid|\operatorname{Gal}(L/K)_{W}|$.
@@
  • Any $W\in\operatorname{Irr}_{L}(H)$ is contained in $V_{L}$ for a unique $V\in\operatorname{Irr}_{K}(H)$. This defines a bijection
    $\operatorname{Irr}_{K}(H)\leftrightarrow\operatorname{Irr}_{L}(H)/\operatorname{Gal}(L/K).$
@@ 10.1. Classical Schur Index
For $W\in\operatorname{Irr}_{L}(H)$, the (classical) Schur index of $W$ over $K$, which we write as $\overline{m}_{K}^{L}(W)$, coincides with the semilinear Schur index for the split extension considered above [HUP, Thm. 38.1]. We see that Proposition 9.3 recovers the following well-known properties of the Schur index.
@@
Let $W\in\operatorname{Irr}_{L}(H)$. Then $W$ is defined over $K$ if and only if $\operatorname{Gal}(L/K)$ fixes $W$ and $\overline{m}_{K}^{L}(W)=1$.
@@
Suppose that $W\in\operatorname{Irr}_{L}(H)$, and set $\mathcal{O}(W)\coloneqq\bigoplus_{\gamma\in\operatorname{Gal}(L/K)/\operatorname{Gal}(L/K)_{W}}\gamma*W$. Then:
@@
    $\mathcal{O}(W)^{\oplus\overline{m}_{K}^{L}(W)}=V_{L}$ for a unique $V\in\operatorname{Irr}_{K}(H)$,
@@
    $\overline{m}_{K}^{L}(W)$ is the unique $m\geq 1$ such that $\mathcal{O}(W)^{\oplus m}$ is induced from an irreducible object of $\operatorname{\mathbf{Rep}}_{K}(H)$,
@@
    For any semisimple $V\in\operatorname{Irr}_{K}(H)$, the multiplicity of $W$ in $V_{L}$ is a multiple of $\overline{m}_{K}^{L}(W)$.
@@
    $W^{\oplus\overline{m}_{L_{W}}^{L}(W)}=U_{L}$ for a unique $U\in\operatorname{Irr}_{L_{W}}(H)$,
@@
    $\overline{m}_{L_{W}}^{L}(W)$ is the unique $m\geq 1$ such that $W^{\oplus m}$ is induced from an irreducible object of $\operatorname{\mathbf{Rep}}_{L_{W}}(H)$,
@@
    For any semisimple $U\in\operatorname{\mathbf{Rep}}_{L_{W}}(H)$, the multiplicity of $W$ in $U_{L}$ is a multiple of $\overline{m}_{L_{W}}^{L}(W)$.
@@ 10.1.1. Galois Splitting Field
In particular, when $H$ is finite, $L_{W}=K(\chi_{W})$ is normal over $K$, being a subfield of a cyclotomic extension. In particular, from Corollary 10.3,
@@ 11. Base Change for Semilinear Representations
Suppose that $W\in\operatorname{Irr}_{F}(H)$, and that $W_{F^{\prime}}=W_{1}^{n_{1}}\oplus\cdots\oplus W_{r}^{n_{r}}$ is a isotypical direct sum decomposition of $W_{F^{\prime}}$. Suppose that $V\in\operatorname{Irr}_{L}^{\rtimes}(G)$ corresponds to $W$, and $V_{i}\in\operatorname{Irr}_{L^{\prime}}^{\rtimes}(G)$ corresponds to $W_{i}$ for i=1,,ri=1,...,r. Then
@@ 12. Character Theory for Semilinear Representations
Suppose in this section that $G$ is finite, and $|G|$ is coprime to $\operatorname{char}(L)$. In this setting, both $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ and $\operatorname{\mathbf{Rep}}_{L}(H)$ are semisimple by Corollary 2.11. For $V,W\in\operatorname{\mathbf{Rep}}_{L}(H)$, we can consider the inner product of characters
@@
which satisfies
$\langle\chi_{V},\chi_{W}\rangle=\dim_{L}\operatorname{Hom}_{L[H]}(V,W).$
@@
For $\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G), VWV\cong W$ if and only if $\chi_{V}=\chi_{W}$.
@@
For $V,W\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$, $\langle\chi_{V},\chi_{W}\rangle=\dim_{K}\operatorname{Hom}_{L\rtimes G}(V,W)$.
@@
In particular, we see that the action of $\Gamma$ on $\operatorname{Fun}(H/\!/H,L)$ preserves the subset $\operatorname{Char}(H,L)$ of characters of objects of $\operatorname{\mathbf{Rep}}_{L}(H)$, and the smaller subset $\operatorname{Irr}(H,L)$ of characters of of irreducible representations of $H$ over $L$. Note that $\operatorname{Irr}(H,L)$ forms a basis for $\operatorname{Fun}(H/\!/H,L)$ whenever $L$ splits $H$, but need not span $\operatorname{Fun}(H/\!/H,L)$ in general.
@@
$\langle\chi_{V},\chi_{W}\rangle=a_{1}b_{1}\dim_{K}D_{1}+\cdots+a_{r}b_{r}\dim_{K}D_{r},$
@@
$\dim_{K}D_{k}=\langle\psi_{k},\psi_{k}\rangle=m_{K}^{L}(\chi_{k})^{2}\cdot|{\mathcal{C}}_{k}|\cdot\langle\chi_{k},\chi_{k}\rangle.$
@@
From Corollary 9.4, we can go further is when $K$ is finite.
@@
Suppose that $K$ is finite. Then $m_{K}^{L}(\chi_{k})=1$ for all k=1,,rk=1,...,r.
In particular, this completely describes the category $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ when $K$ is finite.
@@ 13. Example: $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(S_{3})$ for $L/K$ of Degree $2$ Revisited
In this section we give an example of the theory of the previous sections, to describe the semilinear representations of $S_{3}$ acting on a degree two extension $L/K$ through the quotient by $H\coloneqq\langle(123)\rangle$. We will see that this recovers the classification of Section 5 deduced by more hands-on methods.
@@ 13.1. Characteristic $\neq 3$
Let us first suppose that $\text{char}(K)\neq 3$. From Theorem 8.20 we need to understand the character table of $H$ over $L$, and its action of $\Gamma=\operatorname{Gal}(L/K)$. There are three possible cases, which exactly match and explain the three cases of Section 5.
@@
Let $U$ denotes the irreducible $K$-linear representation of $S_{3}$ with character $\psi_{3}$, and suppose that
@@
$\chi_{\omega}+\chi_{\omega^{2}}=\psi|_{H}=a_{1}m(V_{1})\chi_{1}+a_{\omega}m(V_{\omega})\chi_{\omega}+a_{\omega^{2}}m(V_{\omega^{2}})\chi_{\omega^{2}}.$

### v3

@@ 9. Semilinear Schur Index
For $W\in\operatorname{Ind}_{L}(H)$, set $m_{K}^{L}(W)\coloneqq m(V)$ for the unique corresponding $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$.
@@
Let $W\in\operatorname{Ind}_{L}(H)$. Then $W$ is extendable to a semilinear representation of $G$ if and only if $\Gamma$ fixes $W$ and $m_{K}^{L}(W)=1$.
From the results of the previous section, the numbers $m_{K}^{L}(W)$ have the following properties. Recall that $G_{W}$ is the preimage of $\Gamma_{W}$ in $G$, and we write $L_{W}\coloneqq L^{\Gamma_{W}}$. For any intermediate field $E$ with $K\leq E\leq L$, we write $\Gamma_{E}$ for the preimage of $\operatorname{Gal}(L/E)$ under the isomorphism $\Gamma\xrightarrow{\sim}\operatorname{Gal}(L/K)$.
@@
Suppose that $W\in\operatorname{Ind}_{L}(H)$, and set $\mathcal{O}(W)\coloneqq\bigoplus_{\gamma\in\Gamma/\Gamma_{W}}\gamma*W$. Then:
@@
    $\mathcal{O}(W)^{\oplus m_{K}^{L}(W)}=V|_{H}$ for a unique $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$,
@@
    $m_{K}^{L}(W)$ is the unique $m\geq 1$ such that $\mathcal{O}(W)^{\oplus m}$ extends to an indecomposable object of $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$,
@@
    For any $V\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$, the number of times $W$ appears as a direct summand in the Krull-Remak-Schmidt decomposition of $V|_{H}$ is a multiple of $m_{K}^{L}(W)$.
@@
    $W^{\oplus m_{L_{W}}^{L}(W)}=U|_{H}$ for a unique $U\in\operatorname{Ind}_{L}^{\rtimes}(G_{W})$,
@@
    $m_{L_{W}}^{L}(W)$ is the unique $m\geq 1$ such that $W^{\oplus m}$ extends to an indecomposable object of $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G_{W})$,
@@
    For any $U\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G_{W})$, the number of times $W$ appears as a direct summand in the Krull-Remak-Schmidt decomposition of $U|_{H}$ is a multiple of $m_{L_{W}}^{L}(W)$.
@@
Points (1), (2), (3), (4), (5) and (6) follow from Theorem 8.20 and Corollary 8.6. For point (7), we may take $S$ to be $G_{W}$ in Proposition 8.24, to obtain that $m_{L_{W}}^{L}(W)\mid m_{K}^{L}(W)$ and $m_{K}^{L}(W)\mid m_{L_{W}}^{L}(W)\cdot[\Gamma_{W}:(\Gamma_{W})_{W}]=m_{L_{W}}^{L}(W)$.
@@
with kernel $C_{n}=\langle x\rangle\subset C_{np}=\langle y\rangle$, where $y^{p}=x$. Then for any $1$-dimensional representation $W$ of $H$ valued in $K$, then similarly to above, $m_{K}^{L}(W)\in\{1,p\}$, and $m_{K}^{L}(W)=1$ if and only if $N_{K}^{L}(a)=\chi_{W}(x)$ for some $a\in L^{\times}$.
@@ 10. Recovery of the Classical Schur Index
Let $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$ and let $W\in\operatorname{Ind}_{L}(H)$ be an indecomposable direct summand of $V_{L}$. Then:
@@
    for some integer $\overline{m}_{K}^{L}(W)\geq 1$. In particular, $\overline{m}_{K}^{L}(W)\mid|\operatorname{Gal}(L/K)_{W}|$.
  • $W$ is irreducible if and only if $V$ is irreducible.
@@
  • Any $W\in\operatorname{Ind}_{L}(H)$ is a direct summand of $V_{L}$ for a unique $V\in\operatorname{Ind}_{K}(H)$. Sending $W$ to this unique $V$ defines bijections
    $\operatorname{Ind}_{K}(H)\leftrightarrow\operatorname{Ind}_{L}(H)/\operatorname{Gal}(L/K),\qquad\operatorname{Irr}_{K}(H)\leftrightarrow\operatorname{Irr}_{L}(H)/\operatorname{Gal}(L/K).$
@@ 10.1. Classical Schur Index
For $L/K$ a finite Galois extension and $W\in\operatorname{Ind}_{L}(H)$, we call the number $\overline{m}_{K}^{L}(W)$ appearing in Corollary 10.1 the (classical) Schur index of $W$ over $K$. When $W$ is irreducible (which is the setting in which Schur indices are usually considered) this agrees by definition with the usual definition of the Schur index [HUP, Thm. 38.1].
We following record how Proposition 9.3 recovers the basic properties of the Schur index, and further generalises them to indecomposable modules.
@@
Let $W\in\operatorname{Ind}_{L}(H)$. Then $W$ is defined over $K$ if and only if $\operatorname{Gal}(L/K)$ fixes $W$ and $\overline{m}_{K}^{L}(W)=1$.
@@
Suppose that $W\in\operatorname{Ind}_{L}(H)$, and set $\mathcal{O}(W)\coloneqq\bigoplus_{\gamma\in\operatorname{Gal}(L/K)/\operatorname{Gal}(L/K)_{W}}\gamma*W$. Then:
@@
    $\mathcal{O}(W)^{\oplus\overline{m}_{K}^{L}(W)}=V_{L}$ for a unique $V\in\operatorname{Ind}_{K}(H)$,
@@
    $\overline{m}_{K}^{L}(W)$ is the unique $m\geq 1$ such that $\mathcal{O}(W)^{\oplus m}$ is induced from an indecomposable object of $\operatorname{\mathbf{Rep}}_{K}(H)$,
@@
    For any $V\in\operatorname{Irr}_{K}(H)$, the number of times $W$ appears as a direct summand in the Krull-Remak-Schmidt decomposition of $V_{L}$ is a multiple of $\overline{m}_{K}^{L}(W)$.
@@
    $W^{\oplus\overline{m}_{L_{W}}^{L}(W)}=U_{L}$ for a unique $U\in\operatorname{Ind}_{L_{W}}(H)$,
@@
    $\overline{m}_{L_{W}}^{L}(W)$ is the unique $m\geq 1$ such that $W^{\oplus m}$ is induced from an indecomposable object of $\operatorname{\mathbf{Rep}}_{L_{W}}(H)$,
@@
    For any $U\in\operatorname{\mathbf{Rep}}_{L_{W}}(H)$, the number of times $W$ appears as a direct summand in the Krull-Remak-Schmidt decomposition of $U_{L}$ is a multiple of $\overline{m}_{L_{W}}^{L}(W)$.
@@ 10.1.1. Galois Splitting Field
In particular, when $H$ is finite and $W\in\operatorname{Irr}_{L}(H)$, $L_{W}=K(\chi_{W})$ is normal over $K$, being a subfield of a cyclotomic extension. In particular, from Corollary 10.3,
@@ 11. Base Change for Semilinear Representations
Suppose that $W\in\operatorname{Ind}_{F}(H)$, and that $W_{F^{\prime}}=W_{1}^{n_{1}}\oplus\cdots\oplus W_{r}^{n_{r}}$ is a direct sum decomposition of $W_{F^{\prime}}$, where $W_{i}\in\operatorname{Ind}_{F^{\prime}}(H)$ are pairwise non-isomorphic. Suppose that $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$ corresponds to $W$, and $V_{i}\in\operatorname{Ind}_{L^{\prime}}^{\rtimes}(G)$ corresponds to $W_{i}$ for i=1,,ri=1,...,r. Then
@@ 12. Character Theory for Semilinear Representations
Suppose in this section that $G$ is finite, and $|H|\in L^{\times}$, which by Corollary 8.21 is equivalent to the claim that $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ is semisimple, or that $\operatorname{\mathbf{Rep}}_{L}(H)$ is semisimple. For $V,W\in\operatorname{\mathbf{Rep}}_{L}(H)$, we can consider the inner product of characters
@@
which satisfies [LOR, Lem. 3.20(b)]
$\langle\chi_{V},\chi_{W}\rangle=\dim_{L}\operatorname{Hom}_{L[H]}(V,W)\cdot 1_{L},$
where the multiplication by $1_{L}$ is to ensure this is interpreted as an equality in $L$.
@@
For $V,W\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$, $\langle\chi_{V},\chi_{W}\rangle=\dim_{K}\operatorname{Hom}_{L\rtimes G}(V,W)\cdot 1_{L}$.
@@
If $V,W\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ and $\operatorname{char}(L)=0$, then $V\cong W$ if and only if $\chi_{V}=\chi_{W}$.
@@
In particular, we see that the action of $\Gamma$ on $\operatorname{Fun}(H/\!/H,L)$ preserves the subset $\operatorname{Char}(H,L)$ of characters of objects of $\operatorname{\mathbf{Rep}}_{L}(H)$, and the smaller subset $\operatorname{Irr}(H,L)$ of characters of of irreducible representations of $H$ over $L$. Note that $\operatorname{Irr}(H,L)$ forms a basis for $\operatorname{Fun}(H/\!/H,L)$ whenever $L$ splits $H$ [LOR, Cor. 3.21], but need not span $\operatorname{Fun}(H/\!/H,L)$ in general. However, it is still true that the character of any irreducible representation of $H$ over $L$ is non-zero, irreducible representations are determined by their characters, and that the set $\operatorname{Irr}(H,L)$ is $L$-linearly independent [ISA, Cor. 9.22].
@@
$\langle\chi_{V},\chi_{W}\rangle=(a_{1}b_{1}\dim_{K}D_{1}+\cdots+a_{r}b_{r}\dim_{K}D_{r})\cdot 1_{L},$
@@
$\dim_{K}D_{k}\cdot 1_{L}=\langle\psi_{k},\psi_{k}\rangle=m_{K}^{L}(\chi_{k})^{2}\cdot|{\mathcal{C}}_{k}|\cdot\langle\chi_{k},\chi_{k}\rangle.$
@@
From Corollary 9.4, we can go further is when $L$ is finite.
@@
Suppose that $L$ is finite. Then $m_{K}^{L}(\chi_{k})=1$ for all k=1,,rk=1,...,r.
In particular, the $\Gamma$-set $\operatorname{Irr}(H,L)$ completely describes the category $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ when $L$ is finite.
Corollary 12.8.
Suppose that $W\in\operatorname{Irr}_{L}^{\rtimes}(G)$. Then $\chi_{W}\neq 0$ if and only if $m(W)\cdot 1_{L}\neq 0$, and those characters $\chi_{W}$ where $\chi_{W}\neq 0$ are distinct and $L$-linearly independent. In particular, the association
$\operatorname{Irr}_{L}^{\rtimes}(G)\rightarrow\operatorname{Fun}(H/\!/H,L),\qquad W\mapsto\chi_{W},$
is injective if $L$ has characteristic 0 or if $L$ is finite.
Proof.
This follows directly from Theorem 12.6, Theorem 12.7, and the fact that $\chi_{V}$ for $V\in\operatorname{Irr}_{L}(H)$ are all non-zero, distinct, and $L$-linearly independent [ISA, Cor. 9.22]. ∎
@@ 13. Example: $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(S_{3})$ for $L/K$ of Degree $2$ Revisited
In this section we give an example of the theory of the previous sections, to describe $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(S_{3})$ for any field $L$ where $S_{3}$ acts on $L$ with kernel $H\coloneqq\langle(123)\rangle$ through an order two automorphism, so that $L/K$ has degree two. We will see that this recovers the classification of Section 5 in characteristic 0 deduced by more hands-on methods.
@@ 13.1. Characteristic $\neq 3$
Let us first suppose that $\text{char}(K)\neq 3$, so $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(S_{3})$ is semisimple by Corollary 8.21 and we may use the character theory of Section 12. From Theorem 8.20 we need to understand the character table of $H$ over $L$, and its action of $\Gamma=\operatorname{Gal}(L/K)$. There are three possible cases, which exactly match and explain the three cases encountered in Section 5.
@@
Let $U$ denote the irreducible $K$-linear representation of $S_{3}$ with character $\psi_{3}$, and suppose that
@@
$\chi_{\omega}+\chi_{\omega^{2}}=\psi_{3}|_{H}=a_{1}m(V_{1})\chi_{1}+a_{\omega}m(V_{\omega})\chi_{\omega}+a_{\omega^{2}}m(V_{\omega^{2}})\chi_{\omega^{2}}.$